# Vulnerable Connectivity Caused by Local Communities in Spatial Networks

### v2

@@ Acknowledgements
S26 Figure: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S27 Figure: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S28 Figure: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S29 Figure: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S30 Figure: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S31 Figure: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S32 Figure: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S33 Figure: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S34 Figure: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S35 Figure: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
.6 R to Q scatters
@@
S36 Figure: Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 100 nodes
@@
S37 Figure: Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 10000 nodes
.7 tables
@@
S1 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=100$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. For $R^{RB}$, values with upper-triangles ($\triangle$) indicate higher robustness than Uni. case, while values with lower-triangles ($\triangledown$) indicate lower robustness than Uni. case. For $q_{c}^{RB}$, all values are marked with lower-triangles ($\triangledown$) as they show lower robustness than Uni. case.
@@
S2 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=1024$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. Values with lower-triangles ($\triangledown$) indicate where the cases of Pop. and Inv. have lower robustness of connectivity than the cases of Uni. for both RNG and GG.
@@
S3 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=10000$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. Values with lower-triangles ($\triangledown$) indicate where the cases of Pop. and Inv. have lower robustness of connectivity than the cases of Uni. for both RNG and GG.
@@
S4 Table: Average degree $\langle k\rangle$ in networks with (N) = 100 nodes for seven major Japanese areas. Higher average degrees mean more links per node in the network. Values with upper-triangles ($\triangle$) indicate where the cases of Pop. and Inv. have higher $<k>$ than the cases of Uni. for both RNG and GG.
@@
S5 Table: Average degree $\langle k\rangle$ in networks with (N) = 10000 nodes for seven major Japanese areas. Higher average degrees mean more links per node in the network. Values with upper-triangles ($\triangle$) indicate where the cases of Pop. and Inv. have higher $<k>$ than the cases of Uni. for both RNG and GG.
@@
S6 Table: Modularity $Q$ in networks with 100 nodes for seven major Japanese areas. Higher values indicate stronger community structures. Values with upper-triangles ($\triangle$) or lower-triangles ($\triangledown$) indicate where the cases of Pop. and Inv. have higher or lower modularity than the case of Uni. for both RNG and GG. Note the generally higher modularity in Pop. and Inv. compared to Uni. networks for both RNG and GG.
@@
S7 Table: Modularity $Q$ in networks with 10000 nodes for seven major Japanese areas. Higher values indicate stronger community structures. Values with upper-triangles ($\triangle$) indicate where the cases of Pop. and Inv. have higher modularity than the case of Uni. for both RNG and GG. Note the generally higher modularity in Pop. and Inv. compared to Uni. networks for both RNG and GG.

### v4

@@ Acknowledgements
S26 Fig: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S27 Fig: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S28 Fig: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=1024$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S29 Fig: Robustness against recalculated betweenness (RB) attacks for Fukuoka networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S30 Fig: Robustness against recalculated betweenness (RB) attacks for Hiroshima networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S31 Fig: Robustness against recalculated betweenness (RB) attacks for Keihan networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S32 Fig: Robustness against recalculated betweenness (RB) attacks for Nagoya networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S33 Fig: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S34 Fig: Robustness against recalculated betweenness (RB) attacks for Sendai networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S35 Fig: Robustness against recalculated betweenness (RB) attacks for Sapporo networks with $N=10000$ nodes. For both Rew (Randomized networks) and 2DL lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) $S^{1st}(q)/N$ the relative size of largest connected component, and (c) (d) $S^{2nd}(q)/N$ the critical fraction $q_{c}$ at the peak of the relative size of second largest component.
@@
S36 Fig: Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 100 nodes
@@
S37 Fig: Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 10000 nodes
S38 Fig: Semi-logarithmic plot of population per $500\mathrm{m}\times 500\mathrm{m}$ block meshes in seven major Japanese areas, with meshes sorted in decreasing order of population. Each curve represents an area. The linear decay on the logarithmic scale indicates that a small number of meshes concentrate the majority of the urban population. This heavy-tailed distribution supports the use of rank-based node selections for the Pop. and Inv. networks.
@@
S1 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=100$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. For $R^{RB}$, values with upper-triangles ($\triangle$) indicate higher robustness than Uni.-based case, while values with lower-triangles ($\triangledown$) indicate lower robustness than Uni.-based case. For $q_{c}^{RB}$, all values are marked with lower-triangles ($\triangledown$) as they show lower robustness than Uni.-based case.
@@
S2 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=1024$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. Values with lower-triangles ($\triangledown$) indicate where the cases of Pop.-based and Inv.-based networks have lower robustness of connectivity than the cases of Uni.-based networks for both RNG and GG.
@@
S3 Table: Robustness index ($R^{RB}$) and Critical fraction ($q_{c}^{RB}$) against Recalculated Betweenness (RB) attacks in networks with $N=10000$ nodes for seven major Japanese areas. Higher values indicate greater robustness of connectivity. Values with lower-triangles ($\triangledown$) indicate where the cases of Pop.-based and Inv.-based have lower robustness of connectivity than the cases of Uni.-based for both RNG and GG.
@@
S4 Table: Average degree $\langle k\rangle$ in networks with (N) = 100 nodes for seven major Japanese areas. Higher average degrees mean more links per node in the network. Values with upper-triangles ($\triangle$) indicate where the cases of Pop.-based and Inv.-based networks have higher $<k>$ than the cases of Uni.-based for both RNG and GG.
@@
S5 Table: Average degree $\langle k\rangle$ in networks with (N) = 10000 nodes for seven major Japanese areas. Higher average degrees mean more links per node in the network. Values with upper-triangles ($\triangle$) indicate where the cases of Pop.-based and Inv.-based networks have higher $<k>$ than the cases of Uni. for both RNG and GG.
@@
S6 Table: Modularity $Q$ in networks with 100 nodes for seven major Japanese areas. Higher values indicate stronger community structures. Values with upper-triangles ($\triangle$) or lower-triangles ($\triangledown$) indicate where the cases of Pop.-based and Inv.-based networks have higher or lower modularity than the case of Uni.-based networks for both RNG and GG. Note the generally higher modularity in Pop.-based and Inv.-based networks compared to Uni.-based networks for both RNG and GG.
@@
S7 Table: Modularity $Q$ in networks with 10000 nodes for seven major Japanese areas. Higher values indicate stronger community structures. Values with upper-triangles ($\triangle$) indicate where the cases of Pop.-based and Inv.-based networks have higher modularity than the case of Uni.-based networks for both RNG and GG. Note the generally higher modularity in Pop.-based and Inv.-based networks compared to Uni.-based networks for both RNG and GG.
Network Metric F-value p-value $\eta^{2}$ Distributional Differences
RNG $R^{RB}$ 37.86 $3.56\times 10^{-7}$ 0.808 Uni. $>$ Inv. = Pop.
GG $R^{RB}$ 40.60 $2.13\times 10^{-7}$ 0.819 Uni. $>$ Inv. $>$ Pop.
RNG ${q_{c}}^{RB}$ 98.65 $2.00\times 10^{-10}$ 0.916 Uni. $>$ Inv. $\geq$ Pop.
GG ${q_{c}}^{RB}$ 43.03 $1.39\times 10^{-7}$ 0.827 Uni. $>$ Inv. = Pop.
S8 Table: One-way ANOVA results evaluating the influence of node distribution (Population-based, Inverse population-based, and Uniform) on the robustness index ($R$) and the critical fraction ($q_{c}$) against Recalculate Betweenness (RB) attacks ($N=1024$ nodes). F-value indicates the ratio of between-group variance to within-group variance. p-value shows the probability that the observed group differences are due to chance (typically, $p<0.05$ is considered statistically significant). $\eta^{2}$ denotes the effect size, indicating the proportion of variance explained by the group factor; conventionally, $\eta^{2}>0.14$ is considered a large effect.
Network Metric F-value p-value $\eta^{2}$ Distributional Differences
RNG $R^{ID}$ 13.25 $0.00029$ 0.595 Uni. is significantly stronger than Pop. and Inv.
GG $R^{ID}$ 6.75 $0.0065$ 0.429 Uni. $>$ Inv. $>$ Pop.
RNG ${q_{c}}^{ID}$ 1.50 0.249 0.143 Not significant; inconsistent rankings across cities
GG ${q_{c}}^{ID}$ 15.76 $0.00011$ 0.637 Uni. is significantly stronger than Pop. and Inv.
S9 Table: One-way ANOVA results evaluating the influence of node’s locations (Population-based, Inverse population-based, and Uniform) on the robustness index ($R$) and the critical fraction ($q_{c}$) against Initial Degree (ID) attacks ($N=1024$ nodes). F-value indicates the ratio of between-group variance to within-group variance. p-value shows the probability that the observed group differences are due to chance (typically, $p<0.05$ is considered statistically significant). $\eta^{2}$ denotes the effect size, indicating the proportion of variance explained by the group factor. $\eta^{2}>0.14$ is considered a large effect. In this table, the $q_{c}$ result for networks in RNG shows no significant difference ($p=0.249$). This may be due to two factors: (1) the higher average degree $\langle k\rangle$ of Pop.-based networks in RNG (see Table 4), which offers more alternative paths after node removal; and (2) the presence of grid-like substructures in Pop.-based networks, which can enhance local community and delay critical fragmentation.
Network Metric F-value p-value $\eta^{2}$ Distributional Differences
RNG $R^{RF}$ 2.58 0.104 0.223 Not significant; Pop. and Uni. each ranked highest in 3 cities
GG $R^{RF}$ 22.62 $0.000012$ 0.715 Uni. $>$ Inv. $>$ Pop.
RNG ${q_{c}}^{RF}$ 4.39 $0.028$ 0.328 Uni. is significantly stronger than Pop. and Inv.
GG ${q_{c}}^{RF}$ 5.87 $0.0109$ 0.395 Uni. is significantly stronger than Pop. and Inv.
S10 Table: One-way ANOVA results evaluating the influence of node’s locations (Population-based, Inverse population-based, and Uniform) on the robustness index ($R$) and the critical fraction ($q_{c}$) against Random Failures (RF) ($N=1024$ nodes). F-value indicates the ratio of between-group variance to within-group variance. p-value shows the probability that the observed group differences are due to chance (typically, $p<0.05$ is considered statistically significant). $\eta^{2}$ denotes the effect size, indicating the proportion of variance explained by the group factor; conventionally, $\eta^{2}>0.14$ is considered a large effect. Notably, the $R$ result for networks in RNG shows no significant difference ($p=0.104$), and the $q_{c}$ result is marginally significant ($p=0.028$). These results may be explained by two factors: (1) higher average degree $\langle k\rangle$, which provides alternative paths after node removal (see Table 4); and (2) the presence of grid-like substructures in certain area of the network, which enhance local community and delay critical fragmentation.
S11 Table: Number of detected communities in networks with 1024 nodes for seven major Japanese areas under Original and 2D Lattice conditions
Cities Original 2D Lattice
RNG GG RNG GG
Inv. Pop. Inv. Pop. Inv. Pop. Inv. Pop.
Fukuoka 25 24 21 23 25 24 17 19
Hiroshima 22 27 20 22 28 23 17 20
Keihan 22 23 19 19 27 22 16 19
Nagoya 24 22 20 20 28 23 16 20
Tokyo 23 24 19 21 26 22 16 21
Sendai 23 25 21 22 25 23 20 19
Sapporo 25 22 22 20 25 20 19 22
Uniform 22 20 25 16
Estimated number of communities in original networks and after node relocation on a 2D lattice (relocated). Values with upper-triangles ($\triangle$) indicate that Pop.-based or Inv.-based has more communities than Uni.-based networks. Values in bold indicate fewer communities than Uni.-based networks in both RNG and GG.
Cities RNG GG
Inv. Pop. Inv. Pop.
Fukuoka 0.4235 0.9489 0.4214 0.9437
Hiroshima 0.4163 0.9470 0.4099 0.9405
Keihan 0.4634 0.8842 0.4572 0.8406
Nagoya 0.4638 0.8651 0.4500 0.8254
Tokyo 0.4270 0.8825 0.4181 0.8652
Sendai 0.4660 0.9559 0.4594 0.9468
Sapporo 0.5217 0.9673 0.5139 0.9602
Uniform 0.3025 0.3001
2D Lattice 0.9738 0.9200
S12 Table: Sparsity Index ($SI(G_{w})$) of RNG and GG for different Japanese cities ($N=1024$). Values are calculated based on normalized edge distances, representing the spatial sparsity of each configuration. Uniform and 2D Lattice cases serve as comparative baselines for sparsity. 2D Lattice has the highest $SI(G_{w})$, which may result from long edges between remote nodes.
Measures RB ID RF
RNG GG RNG GG RNG GG
Pearson $r$ ($R$) -0.4071 -0.6514 +0.7017 -0.3532 +0.0786 -0.7300
Pearson $r$ ($q_{c}$) -0.5468 -0.3181 +0.3498 -0.5543 -0.2504 -0.4458
$p$-value ($R$) 0.1486 0.0116 0.0052 0.2155 0.7894 0.0030
$p$-value ($q_{c}$) 0.0430 0.2678 0.2202 0.0397 0.3879 0.1101
S13 Table: Pearson correlation coefficients ($r$) and significance levels ($p$-values) between the sparsity index (SI) and robustness measures ($R$ and $q_{c}$) in RNG and GG against three attack strategies (RB, ID, and RF). Bold values indicate statistically significant results ($p<0.05$). A negative correlation ($r<0$) suggests that higher sparsity (larger SI) is associated with lower robustness, implying that spatially sparser networks are more vulnerable. Conversely, a positive correlation ($r>0$) indicates that sparsity is associated with higher robustness, which may occur in certain cases where redundant links or grid-like structures compensates for sparsity.
Metric RNG GG
$r$ 0.2694 0.5930
$p$-value 0.3315 0.0198
S14 Table: Pearson correlation coefficient ($r$) and significance level ($p$-value) between modularity $Q$ and sparsity index SI($G_{w}$) for RNG and GG. The correlation coefficient $r>0$ indicating a monotonic increasing relation between modularity $Q$ and spatial sparsity $SI(G_{w})$ in RNG and GG, while significant in GG ($p<0.05$).